OpenStudy (anonymous):

The formula for degrees Celsius is: C = 9 over 5 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 8 degrees Celsius is.

OpenStudy (anonymous):

Part 1: C=9/5(F-32) 5C=9(F-32) 5C/9=F-32 F=(5C/9)+32 Part 2: F=[(5x8)/9]+32 F=36.44